# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ 11.3 Passing to $\mathfrak{E}^{\rm max}$
Since $\mathcal{O}_{\rm max}[p^{-1}]\subset B_{\rm max}^{G_{K}}$ has cokernel annihilated by $\varphi^{3}$ by [Breuil, Prop. 5.1.1], we deduce (once again by Remark 11.5) that
@@ 11.6 End of proof of Theorem 3
is an isomorphism. First notice that $\mathcal{L}^{\vee}$ is crystalline associated to the dual of $\bigl(\mathcal{E},\mathrm{Fil}^{\bullet}\mathcal{E},\nabla,\mathfrak{E},\varphi_{\mathfrak{E}}\bigr)$ by Proposition 8.3. We then consider the diagrams of Corollary 11.1 for the cohomologies $\mathrm{H}^{i}$ with compact support of $\mathcal{L}(r)$ and $\bigl(\mathcal{E},\mathrm{Fil}^{\bullet}\mathcal{E},\nabla,\mathfrak{E},\varphi_{\mathfrak{E}}\bigr)$ and for $\mathrm{H}^{2d-i}$ of $\mathcal{L}^{\vee}(-r)$ and the dual of $\bigl(\mathcal{E},\mathrm{Fil}^{\bullet}\mathcal{E},\nabla,\mathfrak{E},\varphi_{\mathfrak{E}}\bigr)$. We have perfect Poincaré dualities in the étale and log de Rham settings and they are compatible after tensoring with $B_{\rm dR}$ by [LLZ, Thm. 4.4.1(4)]. This implies, in particular, that $\mathrm{H}^{2d-i}_{\mathrm{et}}\bigl(U_{\overline{K}},\mathcal{L}^{\vee}(-r)\bigr)[p^{-1}]$ is also a semistable representation as it is dual to $\mathrm{H}^{i}_{\mathrm{et},c}\bigl(U_{\overline{K}},\mathcal{L}(r)\bigr)$. Similarly, we have perfect Poincaré duality in the Hyodo-Kato setting by [EY, Prop. 4.6] and it is compatible with log de Rham Poincaré duality thanks to Lemma 15.4 proved below. This implies that $\upsilon$ and the isomorphism
@@ 12 Hochschild-Serre
$\frac{\mathrm{H}_{\mathrm{logdR}}^{i-1}\left(X,\mathcal{E}(-D)\left(r\right)\right)}{\mathrm{Fir}^{r}\mathrm{H}_{\mathrm{logdR}}^{i-1}\left(X,\mathcal{E}(-D)\left(r\right)\right)}\longrightarrow F^{1}\mathrm{H}^{i}\left(\mathbf{Syn}^{c\text{-}\infty}\left(\mathcal{E},\mathfrak{E},r\right)\right)$
@@
The morphisms $\rho_{\mathrm{Syn}}^{i}\left(r\right)$ (resp. $\rho_{\mathrm{Syn},c}^{i}\left(r\right)$) of Theorem 2 map $F^{1}\mathrm{H}^{i}\left(\mathbf{Syn}\left(\mathcal{E},\mathfrak{E},r\right)\right)$ (respectively $F^{1}\mathrm{H}^{i}\left(\mathbf{Syn}^{c\text{-}\infty}\left(\mathcal{E},\mathfrak{E},r\right)\right)$) to $F^{1}\mathrm{H}_{\mathrm{et}}^{i}\left(U,\mathcal{L}\left(r\right)\right)\left[p^{-1}\right]$ (resp. $F^{1}\mathrm{H}_{\mathrm{et},c}^{i}\left(U,\mathcal{L}\left(r\right)\right)\left[p^{-1}\right]$) and they make the following diagram commutative
@@
(resp. for the compactly supported versions). Here, we identify $D_{\mathrm{dR}}\left(\mathrm{H}_{\mathrm{et}}^{i-1}\left(U_{\overline{K}},\mathcal{L}(r)\right)\right)\simeq\mathrm{H}_{\mathrm{logdR}}^{i-1}\left(X,\mathcal{E}\left(r\right)\right)$ (resp. $D_{\mathrm{dR}}\left(\mathrm{H}_{\mathrm{et},c}^{i-1}\left(U_{\overline{K}},\mathcal{L}(r)\right)\right)\simeq\mathrm{H}_{\mathrm{logdR},c}^{i-1}\left(X,\mathcal{E}\left(r\right)\right)$) via the filtered isomorphism of [DLLZb, Thm. 3.2.7(3)] (resp. [LLZ, Thm. 3.1.10]).
@@
coincides with the composite of the natural map $\mathrm{Gr}^{r}D_{\mathrm{dR}}\bigl(\mathrm{H}^{i-1}_{\mathrm{et}}\bigl(U_{\overline{K}},\mathcal{L}(r)\bigr)\bigr)\to\mathbb{K}$ and the map $\mathrm{Gr}^{r}\mathrm{H}_{\mathrm{logdR}}^{i-1}\left(X,\mathcal{E}\left(r\right)\right)\to\mathrm{Gr}^{r}D_{\mathrm{dR}}\bigl(\mathrm{H}^{i-1}_{\mathrm{et}}\bigl(U_{\overline{K}},\mathcal{L}(r)\bigr)\bigr)$, induced by the map of complexes underlying the definition of $\rho_{\mathrm{Syn}}^{i}(r)$. The latter coincides with the isomorphism of [DLLZb, Thm. 3.2.7(3)] ; see the discussion in §11.1. This concludes the proof.
@@ 13 Gysin morphisms for divisors
One proceeds as in the proof of [LLZ, Prop. 4.3.17]. Let $\epsilon\colon X^{\prime}\to X$ and $\iota\colon F\to X$ be the natural morphisms of log-adic spaces. First of all one upgrades the exact sequence on étale sites to an exact sequence on $X_{\mathrm{pket}}$, using [DLLZa, Lemma 4.5.7 & Lemma 4.5.8] :
@@
It follows from the proof of [LLZ, Prop. 4.3.17] that this produces a distinguished triangle :
@@
We can also assume that $\bigl(\mathcal{E},\mathfrak{E}\bigr)$ are free modules possibly after shrinking the charts and then the statement reduces to the case that $\bigl(\mathcal{E},\mathfrak{E}\bigr)=(\mathcal{O}_{\mathcal{X}},\mathcal{O}_{\mathfrak{X}_{\underline{\beta}}})$ using that $\bigl(\mathcal{E},\mathfrak{E}\bigr)$ is crystalline associated to $\mathcal{L}$ (see Definition 9). In this case the compatibility in the de Rham case, and with a similar argument in the crystalline case, is shown in the proof of [LLZ, Prop. 4.3.17]. ∎
@@
We write $\mathbf{Syn}^{c\text{-}\infty,^{\prime}}(\mathcal{E},\mathfrak{E},r)$ for the complex of $\mathbf{Syn}^{\prime}(\mathcal{E}(-D),\mathfrak{E}(-\mathfrak{D}),r)$ as in §6. As in Lemma 13.1 one proves that it sits in an exact sequence of complexes
$0\to\mathbf{Syn}^{c\text{-}\infty}(\mathcal{E},\mathfrak{E},r)\to\mathbf{Syn}^{c\text{-}\infty,^{\prime}}(\mathcal{E},\mathfrak{E},r)\to\mathbf{Syn}^{c\text{-}\infty}(\mathcal{E}|_{\mathcal{F}},\mathfrak{E}|_{\mathfrak{F}},r-1)[-1]\to 0$
@@ 14 Syntomic polynomial complexes
Proceeding similarly we have variants, namely the one with support conditions $\mathbf{Syn}^{c\text{-}\infty}(\mathcal{E},\mathfrak{E},P,r)$ as in Definition 5 and, as in §4.5 and in Definition 6 for $U\subset\mathcal{X}_{k}^{\rm sm}$ open and $Y\subset\mathcal{X}_{k}$ closed and reduced and containing $U$, the rigid syntomic complexes with or without support conditions, namely $\mathbf{Syn}_{{\rm rig},U\subset Y,\alpha_{k}}\bigl(\mathcal{E},\mathfrak{E},P,r\bigr)$ and $\mathbf{Syn}_{{\rm rig},U^{o}\subset Y,\alpha_{k}}^{c}(\mathcal{E},\mathfrak{E},P,r)$ with support on $U^{o}=U\backslash\mathcal{D}_{k}$, as well as those of §6.4. We simply write
@@
Assume that $\mathcal{X}$ is proper of relative dimension $d$. Then for $K_{P,r}=\mathbf{Syn}(\mathcal{E},\mathfrak{E},P,r)$ or $K_{P,r}=\mathbf{Syn}^{c\text{-}\infty}(\mathcal{E},\mathfrak{E},P,r)$ the group $\mathrm{H}^{i}\left(K_{P,r}\right)$ is finite dimensional as an $L$-vector space for every $i$ and it is 0 for $i>2d+2$.
@@ 15 Trace maps
Assume $\mathcal{X}$ is proper and geometrically connected of relative dimension $d$ (see Remark 15.5 below for a slight generalization). Justified by this properness assumption, we will write $\mathbf{Syn}^{c}$ rather than $\mathbf{Syn}^{c\text{-}\infty}$ when considering the analytic syntomic cohomology. Let $f=[K_{0}:\mathbb{Q}_{p}]$ be the degree of the maximal unramified extension of $\mathbb{Q}_{p}$ in $K$. Assume that the $f$-roots of unity $\mu_{f}$ are in $L$. Consider the Hyodo-Kato cohomology with and without support of the structure sheaf
$\mathrm{H}_{\mathrm{HK},c}^{i}:=\mathrm{H}_{\mathrm{HK}}^{i}\left(\left(\mathcal{X}_{k},\alpha_{\mathcal{X}_{k}}\right)/\mathcal{S}_{0}^{0},\mathfrak{O}_{0}\left(-\mathfrak{D}_{0}\right)\right)_{L},\quad\mathrm{H}_{\mathrm{HK}}^{i}:=\mathrm{H}_{\mathrm{HK}}^{i}\left(\left(\mathcal{X}_{k},\alpha_{\mathcal{X}_{k}}\right)/\mathcal{S}_{0}^{0},\mathfrak{O}_{0}\right)_{L}.$
@@
Suppose $P(X)=\prod_{i\in I}\left(1-\alpha_{i}^{-1}X\right)$ in $1+XL[X]$. We say that $\left(P,r\right)$ is admissible for $\mathcal{X}_{k}$ if $r\geq d+1$ and for every $i\in I$ and every $\zeta\in\mu_{f}$ we have (i) $\alpha_{i}\neq p^{d}\zeta$, (ii) $\alpha_{i}\neq p^{d+1}\zeta$ and (iii) $\alpha_{i}\neq\lambda^{-1}p^{d+1}$ for every eigenvalue $\lambda$ of $\Phi_{L}$ on $\mathrm{H}_{\mathrm{HK}}^{1,N=0}$.
We first provide a criterion for Condition (iii) of the Definition.
@@
Assume $X$ arises from a smooth and projective scheme $X^{\rm alg}$ with projective model $\mathcal{X}^{\rm alg}$ having $p$-adic formal completion $\mathcal{X}$. Then the eigenvalues of $\Phi_{L}$ acting on $\mathrm{H}_{\mathrm{HK}}^{1,N=0}$ are algebraic integers of archimedean norm equal to $p^{\frac{1}{2}}$ or $p$. In particular, $\left(P,r\right)$ is admissible for $\mathcal{X}_{k}$ if $r\geq d+1$ and, for every $i\in I$, the element $\alpha_{i}$ is not of the form $p^{d}\zeta$, $p^{d+\frac{1}{2}}\zeta$ or $p^{d+1}\zeta$ for a root of unity $\zeta$.

### v2

@@ 11.3 Passing to $\mathfrak{E}^{\rm max}$
Since $\mathcal{O}_{\rm max}[p^{-1}]\subset B_{\rm max}^{G_{K}}$ has cokernel annihilated by $\varphi^{3}$ by [BRE97, Prop. 5.1.1], we deduce (once again by Remark 11.5) that
@@ 11.6 End of proof of Theorem 3
is an isomorphism. First notice that $\mathcal{L}^{\vee}$ is crystalline associated to the dual of $\bigl(\mathcal{E},\mathrm{Fil}^{\bullet}\mathcal{E},\nabla,\mathfrak{E},\varphi_{\mathfrak{E}}\bigr)$ by Proposition 8.3. We then consider the diagrams of Corollary 11.1 for the cohomologies $\mathrm{H}^{i}$ with compact support of $\mathcal{L}(r)$ and $\bigl(\mathcal{E},\mathrm{Fil}^{\bullet}\mathcal{E},\nabla,\mathfrak{E},\varphi_{\mathfrak{E}}\bigr)$ and for $\mathrm{H}^{2d-i}$ of $\mathcal{L}^{\vee}(-r)$ and the dual of $\bigl(\mathcal{E},\mathrm{Fil}^{\bullet}\mathcal{E},\nabla,\mathfrak{E},\varphi_{\mathfrak{E}}\bigr)$. We have perfect Poincaré dualities in the étale and log de Rham settings and they are compatible after tensoring with $B_{\rm dR}$ by [LLZ23, Thm. 4.4.1(4)]. This implies, in particular, that $\mathrm{H}^{2d-i}_{\mathrm{et}}\bigl(U_{\overline{K}},\mathcal{L}^{\vee}(-r)\bigr)[p^{-1}]$ is also a semistable representation as it is dual to $\mathrm{H}^{i}_{\mathrm{et},c}\bigl(U_{\overline{K}},\mathcal{L}(r)\bigr)$. Similarly, we have perfect Poincaré duality in the Hyodo-Kato setting by [EY20, Prop. 4.6] and it is compatible with log de Rham Poincaré duality thanks to Lemma 15.4 proved below. This implies that $\upsilon$ and the isomorphism
@@ 12 Hochschild-Serre
$\frac{\mathrm{H}_{\mathrm{logdR}}^{i-1}\left(X,\mathcal{E}(-D)\left(r\right)\right)}{\mathrm{Fir}^{r}\mathrm{H}_{\mathrm{logdR}}^{i-1}\left(X,\mathcal{E}(-D)\left(r\right)\right)}\longrightarrow F^{1}\mathrm{H}^{i}\left(\mathbf{Syn}^{c-\infty}\left(\mathcal{E},\mathfrak{E},r\right)\right)$
@@
The morphisms $\rho_{\mathrm{Syn}}^{i}\left(r\right)$ (resp. $\rho_{\mathrm{Syn},c}^{i}\left(r\right)$) of Theorem 2 map $F^{1}\mathrm{H}^{i}\left(\mathbf{Syn}\left(\mathcal{E},\mathfrak{E},r\right)\right)$ (respectively $F^{1}\mathrm{H}^{i}\left(\mathbf{Syn}^{c-\infty}\left(\mathcal{E},\mathfrak{E},r\right)\right)$) to $F^{1}\mathrm{H}_{\mathrm{et}}^{i}\left(U,\mathcal{L}\left(r\right)\right)\left[p^{-1}\right]$ (resp. $F^{1}\mathrm{H}_{\mathrm{et},c}^{i}\left(U,\mathcal{L}\left(r\right)\right)\left[p^{-1}\right]$) and they make the following diagram commutative
@@
(resp. for the compactly supported versions). Here, we identify $D_{\mathrm{dR}}\left(\mathrm{H}_{\mathrm{et}}^{i-1}\left(U_{\overline{K}},\mathcal{L}(r)\right)\right)\simeq\mathrm{H}_{\mathrm{logdR}}^{i-1}\left(X,\mathcal{E}\left(r\right)\right)$ (resp. $D_{\mathrm{dR}}\left(\mathrm{H}_{\mathrm{et},c}^{i-1}\left(U_{\overline{K}},\mathcal{L}(r)\right)\right)\simeq\mathrm{H}_{\mathrm{logdR},c}^{i-1}\left(X,\mathcal{E}\left(r\right)\right)$) via the filtered isomorphism of [DLL+23, Thm. 3.2.7(3)] (resp. [LLZ23, Thm. 3.1.10]).
@@
coincides with the composite of the natural map $\mathrm{Gr}^{r}D_{\mathrm{dR}}\bigl(\mathrm{H}^{i-1}_{\mathrm{et}}\bigl(U_{\overline{K}},\mathcal{L}(r)\bigr)\bigr)\to\mathbb{K}$ and the map $\mathrm{Gr}^{r}\mathrm{H}_{\mathrm{logdR}}^{i-1}\left(X,\mathcal{E}\left(r\right)\right)\to\mathrm{Gr}^{r}D_{\mathrm{dR}}\bigl(\mathrm{H}^{i-1}_{\mathrm{et}}\bigl(U_{\overline{K}},\mathcal{L}(r)\bigr)\bigr)$, induced by the map of complexes underlying the definition of $\rho_{\mathrm{Syn}}^{i}(r)$. The latter coincides with the isomorphism of [DLL+23, Thm. 3.2.7(3)] ; see the discussion in §11.1. This concludes the proof.
@@ 13 Gysin morphisms for divisors
One proceeds as in the proof of [LLZ23, Prop. 4.3.17]. Let $\epsilon\colon X^{\prime}\to X$ and $\iota\colon F\to X$ be the natural morphisms of log-adic spaces. First of all one upgrades the exact sequence on étale sites to an exact sequence on $X_{\mathrm{pket}}$, using [DLL+02, Lemma 4.5.7 & Lemma 4.5.8] :
@@
It follows from the proof of [LLZ23, Prop. 4.3.17] that this produces a distinguished triangle :
@@
We can also assume that $\bigl(\mathcal{E},\mathfrak{E}\bigr)$ are free modules possibly after shrinking the charts and then the statement reduces to the case that $\bigl(\mathcal{E},\mathfrak{E}\bigr)=(\mathcal{O}_{\mathcal{X}},\mathcal{O}_{\mathfrak{X}_{\underline{\beta}}})$ using that $\bigl(\mathcal{E},\mathfrak{E}\bigr)$ is crystalline associated to $\mathcal{L}$ (see Definition 9). In this case the compatibility in the de Rham case, and with a similar argument in the crystalline case, is shown in the proof of [LLZ23, Prop. 4.3.17]. ∎
@@
We write $\mathbf{Syn}^{c-\infty,^{\prime}}(\mathcal{E},\mathfrak{E},r)$ for the complex of $\mathbf{Syn}^{\prime}(\mathcal{E}(-D),\mathfrak{E}(-\mathfrak{D}),r)$ as in §6. As in Lemma 13.1 one proves that it sits in an exact sequence of complexes
$0\to\mathbf{Syn}^{c-\infty}(\mathcal{E},\mathfrak{E},r)\to\mathbf{Syn}^{c-\infty,^{\prime}}(\mathcal{E},\mathfrak{E},r)\to\mathbf{Syn}^{c-\infty}(\mathcal{E}|_{\mathcal{F}},\mathfrak{E}|_{\mathfrak{F}},r-1)[-1]\to 0$
@@ 14 Syntomic polynomial complexes
Proceeding similarly we have variants, namely the one with support conditions $\mathbf{Syn}^{c-\infty}(\mathcal{E},\mathfrak{E},P,r)$ as in Definition 5 and, as in §4.5 and in Definition 6 for $U\subset\mathcal{X}_{k}^{\rm sm}$ open and $Y\subset\mathcal{X}_{k}$ closed and reduced and containing $U$, the rigid syntomic complexes with or without support conditions, namely $\mathbf{Syn}_{{\rm rig},U\subset Y,\alpha_{k}}\bigl(\mathcal{E},\mathfrak{E},P,r\bigr)$ and $\mathbf{Syn}_{{\rm rig},U^{o}\subset Y,\alpha_{k}}^{c}(\mathcal{E},\mathfrak{E},P,r)$ with support on $U^{o}=U\backslash\mathcal{D}_{k}$, as well as those of §6.4. We simply write
@@
Assume that $\mathcal{X}$ is proper of relative dimension $d$. Then for $K_{P,r}=\mathbf{Syn}(\mathcal{E},\mathfrak{E},P,r)$ or $K_{P,r}=\mathbf{Syn}^{c-\infty}(\mathcal{E},\mathfrak{E},P,r)$ the group $\mathrm{H}^{i}\left(K_{P,r}\right)$ is finite dimensional as an $L$-vector space for every $i$ and it is 0 for $i>2d+2$.
@@ 15 Trace maps
Assume $\mathcal{X}$ is proper and geometrically connected of relative dimension $d$. Justified by this properness assumption, we will write $\mathbf{Syn}^{c}$ rather than $\mathbf{Syn}^{c\text{-}\infty}$ when considering the analytic syntomic cohomology. Let $f=[K_{0}:\mathbb{Q}_{p}]$ be the degree of the maximal unramified extension of $\mathbb{Q}_{p}$ in $K$. Assume that the $f$-roots of unity $\mu_{f}$ are in $L$.
Let $\mathbb{I}$ be the unit object $\left(\mathcal{O},\mathfrak{O}\right)$. The aim of this section is to define a $K_{L}$-valued trace map on $\mathrm{H}^{2d+1}\left(\mathbf{Syn}^{c}\left(\mathbb{I},P,r\right)\right)$, assuming that the pair $(P,r)$ is admissible for $\mathcal{X}_{k}$ according to the following
@@
Suppose that $P\left(X\right)=\prod\nolimits_{i\in I}\left(1-\alpha_{i}^{-1}X\right)\in 1+XL\left[X\right]$. We say that $\left(P,r\right)$ is admissible for $\mathcal{X}_{k}$ if $r\geqslant{}d+1$, and $\alpha$ is not equal to $p^{d}\zeta$ or to $p^{d+1}\zeta$ for some root of unity $\zeta$ in $\mu_{f}$.
Denote by
$\mathrm{H}_{\mathrm{logdR},c}^{i}:=\mathrm{H}_{\mathrm{logdR}}^{i}\left(X,\mathcal{O}_{X}\left(-D\right)\right)_{L}$
and
$\mathrm{H}_{\mathrm{HK},c}^{i}:=\mathrm{H}_{\mathrm{HK}}^{i}\left(\left(\mathcal{X}_{k},\alpha_{\mathcal{X}_{k}}\right)/\mathcal{S}_{0}^{0},\mathfrak{O}_{0}\left(-\mathfrak{D}_{0}\right)\right)_{L}$
the compactly supported logarithmic de Rham cohomology and Hyodo–Kato cohomology of the structure sheaf respectively. Similarly set
$\mathrm{H}_{\mathrm{an},c}^{2d}:=\mathrm{H}_{\mathrm{an}}^{2d}\left(\left(\mathcal{X}_{k},\alpha_{\mathcal{X}_{k}}\right)/K_{0},\mathfrak{O}^{\mathrm{conv}}\left(-\mathfrak{D}\right)\right),$
and denote by
$\gamma^{\mathrm{HK}}_{c}:\mathrm{H}_{\mathrm{an},c}^{i}\stackrel{{\scriptstyle}}{{\longrightarrow}}\mathrm{H}^{i}_{\mathrm{HK},c}$
the Frobenius equivariant morphism defined by $Z=0$. According to Corollary 11.1 the map $\gamma_{c}^{\mathrm{an}}:\mathrm{H}^{i}_{\mathrm{an},c}\stackrel{{\scriptstyle}}{{\longrightarrow}}\mathrm{H}^{i}_{\mathrm{logdR},c}$ factors as the composition $\mathrm{comp}_{c}\circ{}\gamma^{\mathrm{HK}}_{c}$, where
$\mathrm{comp}_{c}:\mathrm{H}^{i}_{\mathrm{HK},c}\stackrel{{\scriptstyle}}{{\lhook\joinrel\relbar\joinrel\rightarrow}}\mathrm{H}_{\mathrm{logdR},c}^{i}$
is the inclusion arising from the (compactly supported analogue of the) commutative diagram displayed in the statement of Corollary 11.1. In particular the image of $\gamma_{c}^{\mathrm{an}}$ is contained in the image of the comparison map $\mathrm{comp}_{c}$.
Assume that
$r\geqslant d+1\ \ \text{ and}\ \ P\left(p^{d}\zeta\right)\neq 0$
for any $\zeta\in\mu_{f}$. The Hyodo-Kato trace map identifies $\mathrm{H}_{\mathrm{HK},c}^{2d}$ with $K_{0,L}\left(-d\right)$ (cf. Proposition 4.6 of [EY20]), hence the assumption $P\left(p^{d}\zeta\right)\neq 0$, for $\zeta\in{}\mu_{f}$, implies that $P\left(\Phi_{L}\right)$ is invertible on $\mathrm{H}_{\mathrm{HK},c}^{2d}$. For each $x$ in $\mathrm{H}^{2d}_{\mathrm{an},c}$, we can then define
$P\left(\Phi_{L}\right)^{-1}\left(\gamma_{c}^{\mathrm{an}}\left(x\right)\right)=\mathrm{comp}_{c}\big(P(\Phi_{L})^{-1}\big(\gamma^{\mathrm{HK}}_{c}(x)\big)\big)$
in $\mathrm{H}_{\mathrm{logR},c}^{2d}$. Because $r\geqslant d+1$, one has
$\mathrm{H}^{2d}\big(\mathrm{Gr}\big(\Omega_{X/K}^{\mathrm{log},\bullet}\left(-D\right)\big)\big)_{L}=\mathrm{H}_{\mathrm{logdR},c}^{2d},$
so that
$\mathrm{F}^{1}\mathrm{H}^{2d+1}\left(\mathbf{Syn}^{c}\left(\mathbb{I},P,r\right)\right)=\mathrm{coker}\Big(\mathrm{H}^{2d}_{\mathrm{an},c}\stackrel{{\scriptstyle P(\Phi_{L})\oplus{}\gamma^{\mathrm{an}}_{c}}}{{\longrightarrow}}\mathrm{H}^{2d}_{\mathrm{an,c}}\oplus{}\mathrm{H}^{2d}_{\mathrm{logdR},c}\Big).$ (9)
Since $\gamma_{c}^{\mathrm{HK}}$ is Frobenius equivariant, we can then define
$\partial:\mathrm{F}^{1}\mathrm{H}^{2d+1}\left(\mathbf{Syn}^{c}\left(\mathbb{I},P,r\right)\right)\stackrel{{\scriptstyle}}{{\longrightarrow}}\mathrm{H}^{2d}_{\mathrm{logdR},c}$
to be the map which on the class $[x,y]$ represented by the pair $(x,y)$ takes the value
$\partial\big([x,y]\big)=y-P(\Phi_{L})^{-1}(\gamma^{\mathrm{an}}_{c}(x)).$
Remark 15.1.
Assume that $r\geqslant d+1$ and $P\left(p^{d}\zeta\right)\neq 0$ for any $\zeta\in\mu_{f}$. The above map $\partial$ is a section of the canonical map from $\mathrm{H}_{\mathrm{logR},c}^{2d}$ to $\mathrm{F}^{1}\mathrm{H}^{2d+1}\left(\mathbf{Syn}^{c}\left(\mathbb{I},P,r\right)\right)$ sending $y$ to the class of $\left(0,y\right)$.
The following lemma explains the relevance of the last assumption in Definition 13, namely $P\left(p^{d+1}\zeta\right)\neq 0$ for any $\zeta\in\mu_{f}$.
@@
If $P\left(p^{d+1}\zeta\right)\neq 0$ for any $\zeta\in\mu_{f}$, then
$\mathrm{H}^{2d+1}\left(\mathbf{Syn}^{c}\left(\mathbb{I},P,r\right)\right)=\mathrm{F}^{1}\mathrm{H}^{2d+1}\left(\mathbf{Syn}^{c}\left(\mathbb{I},P,r\right)\right)\text{.}$